# Unlocking True Elasticity for the Cloud-Native Era with Dandelion

### v1

@@ 4.2. Language and Compilation Support
Dandelion provides software development kits (SDK) for functions written in C and C++. These SDKs help users compile their functions against hlibc/hlibc++ or the low-level interface. To support Python functions, we have compiled the CPython interpreter using our C SDK. The same approach can be taken to support other interpreted languages with interpreters written in C or C++. To support more compiled langauges, we aim to extend our SDK to be compatible with LLVM, such that any language with an LLVM frontend can be easily compiled for Dandelion. Additional languages can also be supported by Dandelion’s rWasm backend (§6.2), as it accepts Wasm code as input, which is a compilation target for many languages (emscripten, ).
@@ 4.3. Porting Applications to Dandelion
Dandelion users currently need to express their applications as DAGs of pure compute and communication functions. Some applications like data analytics naturally fit this model, as query engines represent SQL queries as DAGs of operators. We also find that cloud-native applications like image or log processing pipelines that ingest data from cloud storage fit this programming model well. Based on our initial exploration with a Text2SQL agentic application in §7.7, Dandelion may also be suitable for some agentic AI workflows, which are an emerging class of applications that involve interactions between AI inference services, vector databases for retrieval, and custom user logic for pre- and post-processing data between these services.
@@
However, developers may also want to port existing cloud applications originally developed assuming a POSIX interface. For applications that use asynchronous programming (rust-async, ) (supported in popular languages like Python, Rust, and C++), we expect that Dandelion could use a compiler to automatically decompose an application into its pure compute and communication components by dividing the application into a graph of continuations (flanaganEssenceCompilingContinuations1993, ). Continuations are language-level primitives that can save the program state at arbitrary locations (e.g., at communication I/O boundaries) and resume execution at a later time. We plan to explore this approach as future work, drawing inspiration from VectorVisor (ginzburgVectorVisorBinaryTranslation, ), which uses continuations to port code containing system calls to run on GPUs.
@@ 5. Dandelion Execution System
The dispatcher manages memory and compute resources using two abstractions, memory contexts and engines, respectively. The dispatcher prepares an isolated memory context for each ready function and enqueues the function in a compute engine or in a communication engine queue for execution. When an engine finishes executing a function, the dispatcher associates its output data to the input of waiting functions. When a waiting function becomes ready, the dispatcher ensures that the outputs from prior functions are copied as inputs into the new function’s context. The dispatcher de-allocates a completed function’s memory context when all data-dependent functions have consumed its output.
@@ 6.1. Dispatcher and Engine Coordination
Fault tolerance. Pure compute functions are idempotent. If a compute engine reports that a function failed, the dispatcher can re-schedule it. Communication function failures are more complicated due to side effects. Protocol specifications can help Dandelion decide which functions can be re-executed and which functions need special handling (apiary, ; ft-shim, ). For example, HTTP PUT requests are idempotent.
@@ 6.2. Compute Engines
rWasm-based isolation.   Dandelion can also use isolation mechanisms that do not require hardware support. We demonstrate this by implementing a backend using rWasm (rwasm, ), a Wasm-to-safe-Rust compiler. Developers register compute functions in Wasm code. Dandelion transpiles the user-provided Wasm code to safe Rust code, then wraps and compiles it to a native shared library. To execute functions, Dandelion only needs to load the library and call into it using the symbols defined in the wrapper. Memory isolation is provided by the memory safety guarantees for safe code by the Rust compiler, which is already part of the trusted computing base as Dandelion is implemented in Rust. Since memory isolation is enforced at compile time, this approach does not rely on a runtime like Wasmtime for isolation.
@@ 7.1. Methodology
Baselines.   We compare Dandelion to Firecracker (firecracker, ), the open-source MicroVM hypervisor used by AWS Lambda. We use a simple HTTP relay to route requests to MicroVMs and spawn new MicroVMs. When the relay receives a request for a hot function, it passes it to the first available MicroVM. When it gets a cold request, it boots a new MicroVM and forwards the request on as soon as that VM is ready. We use the same setup for gVisor (gvisor, ), with containers instead of MicroVMs. We run Firecracker (which we abbreviate as FC) with and without snapshots (firecracker:snapshot_restoration, ). We also include Spin (spin, ), an open-source framework for building and running event-driven applications using Wasmtime (wasmtime, ). We use Spin’s recommended default settings, which include enabling pooled allocation with 1000 Wasmtime instances. For simplicity, we refer to the Spin/Wasmtime setup as Wasmtime or WT.
@@
Table 1. Dandelion avg. latency breakdown in $\mu$s for each isolation backend running 1x1 matmul on Morello.
@@ 7.6. Multiplexing Applications
In Figure 8, we run the distributed log processing application from Figure 3 (I/O-intensive) and an image compression application that transforms an 18kB QOI (qoi-compression, ) image to PNG (compute-intensive) on Dandelion, Firecracker (with snapshot), and Wasmtime under a bursty load pattern. Consistent with previous experiments, we use a 97% hot request ratio for Firecracker.
Firecracker has a bi-modal latency distribution for both applications. This is because 97% of requests are served by hot VMs, which benefit from pre-warmed state, while the remaining 3% incur cold-start (snapshot restoration) overhead. Wasmtime, on the other hand, struggles to handle the mixed workload effectively – both applications show significantly higher latencies and fluctuations. As discussed in §7.3, Wasmtime suffers from less efficient code generation and optimization. Spin’s reliance on the Rust Tokio async runtime also exacerbates interference between applications (tokio-runtime, ; tokio-scheduling-blog, ). We observe that image compression tasks hog resources, causing log processing tasks to queue behind them, increasing tail latency.
Dandelion maintains stable latency for both applications, even during load bursts, due to fast cold starts and the controller’s dynamic re-allocation of CPU cores, e.g., it scales from a single I/O core at the beginning of the experiment to 4 I/O cores to accommodate I/O-intensive log processing during peak load. For the image compression application, Dandelion achieves an average latency of 18.23 ms, which is 10.6% lower than Firecracker and 65.8% lower than Wasmtime. Its p99 latency of 19.64 ms is 38.6% lower than Firecracker and 67.1% lower than Wasmtime. For the log processing application, Dandelion delivers an average latency of 27.92 ms, which is 9.1% higher than Firecracker (which overprovisions memory to maximize warm starts) and 3.4% lower than Wasmtime. Its p99 latency of 30.61 ms is 19.5% lower than Firecracker and 33.5% lower than Wasmtime. Dandelion reduces performance variability, achieving lowest relative variance of 1.30% for image compression and 2.87% for log processing, compared to Firecracker (389.6% and 1495.17%) and Wasmtime (6.11% and 79.2%). Overall, Dandelion efficiently multiplexes requests with different compute and I/O intensities and maintains low, stable latency despite creating a sandbox on the critical path for each request.
@@ 8. Security Analysis
Trusted computing base.   Dandelion, Firecracker and Spin have several TCB components in common: the Linux kernel, C standard library, and Rust compiler. While the Firecracker and Spin code base has expanded to $\sim$68k and $\sim$65k lines of Rust code respectively, Dandelion consists of $\sim$12k lines of Rust code including tests, with only 2k lines directly relevant to isolation and interaction with user code. gVisor has $\sim$38k lines of Go, excluding thrid-party packages.
@@ 9. Related Work
Serverless programming models. Dandelion draws inspiration from other dataflow paradigms for serverless (compute-centric-networking, ; func_as_func, ; lambdata, ; dataflower, ). Lambdata (lambdata, ) and Dataflower (dataflower, ) leverage dataflow information to reduce function triggering latency and avoid round-trips to remote storage for data exchange. Pheromone (yu2023following, ) exposes a data bucket abstraction (for intermediate function outputs) and trigger primitives, which allows developers to specify when and how to pass intermediate data to other functions and invoke them. Other FaaS systems introduce data dependency hints to optimize function placement across nodes (abdi2023palette, ; mahgoub:sonic, ; faasflow, ). Nu (nu-nsdi, ) provides a new process abstraction that decomposes an application into fine-grained units of state and compute, which can be quickly migrated to improve resource utilization. However, Nu only provides process-level isolation, which is not sufficient for multi-tenant cloud environment.
@@ 10. Conclusion
The cloud has evolved in the past decade to offer a variety of serverless compute, storage, and AI/analytics services, but the performance and efficiency of these services is limited by the legacy software that they retrofit. Today’s infrastructure is still based on the older, more traditional model of cloud computing, in which users rent long-running VMs, each with a guest OS exposing a POSIX-like interface. Initializing this system stack is slow and having applications interact with the platform on the level of files and sockets also does not provide the platform with information like compute vs. I/O intensity to optimize resource allocation and scheduling.
Dandelion co-designs a cloud-native programming model and function execution system for truly elastic computing while maintaining secure isolation. Developers specify their applications as DAGs of pure compute functions and communication functions. We demonstrate how Dandelion can isolate pure compute functions with lightweight sandboxes that boot in 100s of $\mu s$, enabling true elasticity (i.e., fast cold start for every request). Dandelion improves elastic query processing latency by 40% compared to AWS Athena and reduces committed memory by 96% on average compared to Knative autoscaling for the Azure Functions trace.
@@ 11. Acknowledgements
We would like to thank Gustavo Alonso, Timothy Roscoe, Xiaozhe Yao, and Maximilian Böther for their helpful insights during discussions. We would also like to acknowledge Roberto Starc, Lorenzo Rosa, Georgijs Vilums, and Josef Schönberger for their contributions to the project. Thank you to Eric Van Hensbergen for his inputs in the early stages of the project and his support with the Arm Morello setup. We also thank our anonymous reviewers from ASPLOS’25 and OSDI’25 for their feedback and suggestions.

### v2

@@ 4.2. Language and Compilation Support
Dandelion provides software development kits (SDK) for functions written in C and C++. These SDKs help users compile their functions against dlibc/dlibc++ or the low-level interface. To support Python functions, we have compiled the CPython interpreter using our C SDK. The same approach can be taken to support other interpreted languages with interpreters written in C or C++. To support more compiled langauges, we aim to extend our SDK to be compatible with LLVM, such that any language with an LLVM frontend can be easily compiled for Dandelion. Additional languages can also be supported by Dandelion’s rWasm backend (§6.2), as it accepts Wasm code as input, which is a compilation target for many languages (emscripten, ).
@@ 4.3. Porting Applications to Dandelion
Dandelion users currently need to express their applications as DAGs of pure compute and communication functions. Some applications like data analytics naturally fit this model, as query engines represent SQL queries as DAGs of operators. We also find that cloud-native applications like image or log processing pipelines that ingest data from cloud storage fit this programming model well. Based on our initial exploration with a Text2SQL agentic application in §7.7, Dandelion can also be suitable for some agentic AI workflows, which are an emerging class of applications that involve interactions between AI inference services, vector databases for retrieval, and custom user logic for pre- and post-processing data between these services.
1auth_token = format_request(args[1])
2auth_response = http_get(auth_url, auth_token)
3authorized_servers = auth_processing(auth_response)
4responses = []
5for sever in authorized_servers:
6 responses.push(http_get(server))
7rendered_html = <html template start>
8for response in responses:
9 if response.successful():
10 rendered_html.append(reponse.render())
11 else:
12 rendered_html.append(response.error())
13rendered_html.append(<html template end>)
Listing 1: Pseudocode for log processing application.
However, users may also wish to port existing applications to Dandelion. To illustrate this process, consider the pseudocode in Listing 1 for the log processing example application from Figure 8. To port this application, the developer needs to be split the compute logic into separate functions between each of the two HTTP calls that the application issues. This results in three separate compute functions. The first compute function, called Access, contains the parsing logic for the input token and forms the request (line 1 from Listing 1 ). The second compute function, called FanOut, contains the logic from lines 3-5 in Listing 1 while the third compute function, called Render, contains the logic from lines 7-13. The HTTP requests issued on lines 2 and 6 in Listing 1 are done by communication functions in Dandelion. Listing 2 shows the composition that the user would write to express how Dandelion should pass data between compute and communication functions. The each keyword on both of the HTTP functions tells Dandelion it can parallelize the requests, while the all keywords on the compute functions indicate that all results of the preceding functions are collected and fed to a single instance of the function.
1composition RenderLogs(AccessToken) => HTMLOutput {
2 Access(AccessToken = all AccessToken)
3 => (AuthRequest = HTTPRequest);
4 HTTP(Request = each AuthRequest)
5 => (AuthResponse = Response);
6 FanOut(HTTPResponse = all AuthResponse)
7 => (LogRequests = HTTPRequests);
8 HTTP(Request = each LogRequests)
9 => (LogResponses = Response);
10 Render(HTTPResponses = all LogResponses)
11 => (HTMLOutput = HTMLOutput);
12}
Listing 2: Log processing application from Listing 1 expressed as a Dandelion composition DAG using our DSL.
@@ 10 Render(HTTPResponses = all LogResponses)
For applications that use asynchronous programming (rust-async, ) (supported in popular languages like Python, Rust, and C++), we expect that Dandelion could use a compiler to automatically decompose an application into its pure compute and communication components by dividing the application into a graph of continuations (flanaganEssenceCompilingContinuations1993, ). Continuations are language-level primitives that can save the program state at arbitrary locations (e.g., at communication I/O boundaries) and resume execution at a later time. We plan to explore this approach as future work, drawing inspiration from VectorVisor (ginzburgVectorVisorBinaryTranslation, ), which uses continuations to port code containing system calls to run on GPUs.
4.4. Fault handling
When a communication function fails, it forwards the appropriate error code to the next downstream function(s). For example, an HTTP function may return 404 Not Found as a response. Dandelion compositions support conditional control flow, so failures can be handled gracefully. By default, functions are only executed if each of the input sets contains at least one item. This allows certain functions to only run if a failure message is present and to skip execution of further downstream functions if the failure cannot be recovered from. Sets that do not need to contain an item for the function to run can be marked as optional.
Dandelion’s failure handling could benefit from additional techniques, such as the log-based approaches in Boki (jiaBokiStatefulServerless2021, ), Beldi (qiHalfmoonLogOptimalFaultTolerant2023, ) or Halfmoon (qiHalfmoonLogOptimalFaultTolerant2023, ). These systems can be used to find reentry points in compositions and skip functions that have successfully committed their progress.
@@ 5. Dandelion Execution System
The dispatcher manages memory and compute resources using two abstractions: memory contexts and engines, respectively. The dispatcher prepares an isolated memory context for each ready function and enqueues the function in a compute or communication engine queue for execution. When an engine finishes executing a function, the dispatcher associates its output data to the input of waiting functions. When a waiting function becomes ready, the dispatcher ensures that the outputs from prior functions are copied as inputs into the new function’s context. The dispatcher de-allocates a completed function’s memory context when all data-dependent functions have consumed its output.
@@ 6.2. Compute Engines
rWasm-based isolation.  Dandelion can also use isolation mechanisms that do not require hardware support. We show this by implementing a backend using rWasm (rwasm, ), a Wasm-to-safe-Rust compiler. Developers register compute functions in Wasm code. Dandelion transpiles the user-defined Wasm code to safe Rust code, then wraps and compiles it to a native shared library. To execute functions, Dandelion only needs to load the library and call into it using the symbols defined in the wrapper. Memory isolation is provided by the memory safety guarantees for safe code by the Rust compiler, which is already part of the trusted computing base as Dandelion is implemented in Rust. Since memory isolation is enforced at compile time, this approach does not rely on a runtime like Wasmtime for isolation.
@@ 7.1. Methodology
Baselines.  We compare Dandelion to Firecracker (firecracker, ), the open-source MicroVM hypervisor used by AWS Lambda. We use a simple HTTP relay to route requests to MicroVMs and spawn new MicroVMs. When the relay receives a request for a hot function, it passes it to the first available MicroVM. When it gets a cold request, it boots a new MicroVM and forwards the request on as soon as that VM is ready. We run Firecracker (which we abbreviate as FC) with and without snapshots (firecracker:snapshot_restoration, ). The Firecracker baseline represents VM-based isolation approaches. Dandelion’s most similar backend for comparison is the KVM backend, as it relies on the same hardware virtualization.
We use the same HTTP relay setup for gVisor (gvisor, ), which uses containers instead of MicroVMs. The most comparable isolation backend in Dandelion is the process backend, as it also uses processes with restricted capabilities.
We also run Spin (spin, ), an open-source framework for building and running event-driven applications with Wasmtime (wasmtime, ). We use Spin’s recommended default settings, which include enabling pooled allocation with 1000 Wasmtime instances. For simplicity, we refer to the Spin/Wasmtime setup as Wasmtime or WT. This baseline represents software fault isolation and is most suitable to compare to Dandelion’s rWasm isolation backend.
@@
[…]
@@ 7.6. Multiplexing Applications
In Figure 8, we run the distributed log processing application from Figure 3 (I/O-intensive) and an image compression application that transforms an 18kB QOI (qoi-compression, ) image to PNG (compute-intensive) on Dandelion, Firecracker (with snapshot), and Wasmtime under a bursty load pattern. Image compression is as an example of image processing and log processing is an example of HTML templating, which are both common task from FaaS benchmarks (copikSeBSServerlessBenchmark2021, ; kimFunctionBenchSuiteWorkloads2019, ; vSwarm, ). Consistent with previous experiments, we use a 97% hot request ratio for Firecracker.
Firecracker has a bi-modal latency distribution for both applications. This is because 97% of requests are served by hot VMs, which benefit from pre-warmed state, while the remaining 3% incur cold-start (snapshot restoration) overhead. By overprovisioning memory to maximize warm starts, Firecracker achieves a relatively low average request latency for both compression (20.4ms) and log processing (25.6ms), but has high variance at 389% and 1495% respectively due to the cold starts. Spin/Wasmtime, on the other hand, struggles to handle the mixed workload effectively, increasing the latency of both applications. As discussed in §7.3, Wasmtime suffers from less efficient code generation and optimization. Spin’s reliance on the Rust Tokio async runtime also exacerbates interference between applications (tokio-runtime, ; tokio-scheduling-blog, ). Spin schedules functions cooperatively, meaning a thread can be occupied by a compute intensive function for an arbitrary amount of time. We observe that image compression tasks hog resources, causing log processing tasks to queue behind them, increasing tail latency. While this keeps the variance of the compression application relatively low at 6.1%, this is not a good tradeoff as the average latency is 53.3ms. The log processing application has a 79.2% variance with an average latency of 28.9ms.
Dandelion maintains stable latency for both applications, even during load bursts, due to fast cold starts and the controller’s dynamic re-allocation of CPU cores, e.g., it scales from a single I/O core at the beginning of the experiment to 4 I/O cores to accommodate I/O-intensive log processing during peak load. For the image compression application, Dandelion achieves an average latency of 18.2ms and p99 latency of 19.6ms, which is 38.6% lower than Firecracker and 67.1% lower than Wasmtime. For the log processing application, Dandelion delivers an average latency of 27.9ms and p99 latency of 30.6ms, which is 19.5% lower than Firecracker and 33.5% lower than Wasmtime. Dandelion reduces performance variability, achieving the lowest relative variance of 1.3% for image compression and 2.9% for log processing. Overall, Dandelion efficiently multiplexes requests with different compute and I/O intensities and maintains low and stable latency despite creating a sandbox on the critical path for each request.
@@ 8. Security Analysis
Trusted computing base.  Dandelion, Firecracker and Spin have several TCB components in common: the Linux kernel, C standard library, and Rust compiler. While the Firecracker and Spin code base has expanded to $\sim$68k and $\sim$65k lines of Rust code respectively, Dandelion consists of $\sim$12k lines of Rust code including tests, with only 2k lines directly relevant to isolation and interaction with user code. gVisor has $\sim$38k lines of Go, excluding third-party packages.
@@ 9. Related Work
Serverless programming models. Dandelion’s design is closely related to other dataflow paradigms for serverless. Computation-centric networking (compute-centric-networking, ) involves a separation between I/O and compute, with delineated nondeterminism. Lambdata (lambdata, ) and Dataflower (dataflower, ) leverage application dataflow information to reduce function triggering latency and avoid round-trips to remote storage for data exchange. Pheromone (yu2023following, ) exposes a data bucket abstraction (for intermediate function outputs) and trigger primitives, which allows developers to specify when and how to pass intermediate data to other functions and invoke them. Other FaaS systems introduce data dependency hints to optimize function placement across nodes (abdi2023palette, ; mahgoub:sonic, ; faasflow, ). Nu (nu-nsdi, ) provides a new process abstraction that decomposes an application into fine-grained units of state and compute, which can be quickly migrated to improve resource utilization. However, Nu only provides process-level isolation, which is not sufficient for multi-tenant cloud environment.
@@
Data processing systems.  DAG-based systems for data processing such as Spark(spark, ) are already popular. These systems usually focus on single users or groups of mutually trusting users. Dandelion innovates on the DAG model for processing, as it leverages it to achieve secure isolation and high elasticity in multi-tenant environments.
@@ 10. Conclusion
The cloud has evolved in the past decade to offer a variety of serverless compute, storage, and AI/analytics services, but the performance and efficiency of these services is limited by the legacy software that they retrofit. Today’s infrastructure is still based on the older, more traditional model of cloud computing, in which users rent long-running VMs, each with a guest OS exposing a POSIX-like interface. Initializing this system stack is slow. Furthermore, having applications interact with the platform on the level of files and sockets does not provide the platform with useful information about application characteristics, like compute vs. I/O intensity, to optimize resource allocation and scheduling.
Dandelion co-designs a declarative cloud-native programming model and execution system for truly elastic computing while maintaining secure isolation. Developers specify their applications as DAGs of pure compute functions and communication functions. This enables Dandelion to isolate pure compute functions with lightweight sandboxes that boot in 100s of microseconds, enabling true elasticity (i.e., fast cold start for every request). Dandelion improves elastic query processing latency by 40% compared to AWS Athena and reduces committed memory by 96% on average compared to Knative autoscaling for the Azure Functions trace.
@@ 11. Acknowledgements
We would like to thank Gustavo Alonso, Timothy Roscoe, Xiaozhe Yao, and Maximilian Böther for their helpful insights during discussions. We would also like to acknowledge Roberto Starc, Lorenzo Rosa, Georgijs Vilums, and Josef Schönberger, and Alessio Russo for their contributions to the project. Thank you to Eric Van Hensbergen for his inputs in the early stages of the project and his support with the Arm Morello setup. We also thank our anonymous reviewers from ASPLOS’25 and OSDI’25 for their feedback and suggestions. Tom Kuchler, Pinghe Li, and Yazhuo Zhang are supported by the Swiss National Science Foundation (project TMSGI2_218019).